OpenStudy (anonymous):

solve the equation....(x-4)(2x+9)=0

OpenStudy (anonymous):

I came up with the same thing but do they have to be in this order?

OpenStudy (anonymous):

@mwilliamson

OpenStudy (anonymous):

or would it be wrong if I put the 4 first

OpenStudy (anonymous):

i would go with \[ x = 4, - \frac{ 9 }{ 2 }\] but I am not 100% positive on that, see if someone could double check it, what happened to my reply? err this site takes some getting use to,,

hero (hero):

When it says to solve (x - 4)(2x + 9) = 0, what they mean for you to is set each binomial equal to zero, then solve for x: x - 4 = 0 2x + 9 = 0

OpenStudy (anonymous):

mwilliamson is right.

hero (hero):

His answer was right, however he gave the answer (something OS doesn't encourage), so that wasn't right to do.

OpenStudy (anonymous):

@mwilliamson the moderator erased them because they thought I asked the same question twice

OpenStudy (anonymous):

Yeah. I understand. He should have shown how and why he got it.

OpenStudy (anonymous):

well sometimes the asnwer is needed in order for some of us to understand, I know a lot of the time if I can't see the answer, I can't figure out how the problem is worked or need to be worked

OpenStudy (anonymous):

so thanks @mwilliamson, I see no harm

hero (hero):

No, you need to understand how to do the problem. Giving the answer only doesn't help you understand how to do it. It just gives you the answer without you having learned anything @mathgirl501

OpenStudy (anonymous):

and like I said I had the answer, I just need to know if it matters which number is first

OpenStudy (anonymous):

well sometimes seeing the answer is what it takes to understand how to solve it like I said

OpenStudy (anonymous):

i originally put solution 1: \[x = 4\] solution 2:\[x = -\frac{ 9 }{ 2 }\] before they got erased, was that wrong to do it that way?

OpenStudy (anonymous):

That too is also a good point but this is coming from you, "the asker of this question" so do you know how to do this now?

OpenStudy (anonymous):

I never needed help understanding it...you guys aren't listening

OpenStudy (anonymous):

I just need to know one thing

OpenStudy (anonymous):

no @mwilliamson, your fine

hero (hero):

@mwilliamson, no it is not acceptable.

OpenStudy (anonymous):

lol well I guess I dunno, can we get back to the question though

hero (hero):

For those who "don't know" refer to the code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

thanks @Hero, I appreciate all your help with learning this site & with learning Algebra, Ill try to fully explain for now on

OpenStudy (anonymous):

ok great....question.....?????

OpenStudy (anonymous):

please....anyone, I feel like we've turned into facebook

OpenStudy (anonymous):

@mathgirl501 it doesn't matter what order you put them in because they are both "x" solutions.

OpenStudy (anonymous):

Thanks @onelove94

OpenStudy (anonymous):

case closed guys

hero (hero):

@mathgirl501, does facebook help you with your homework? If it does, why come here?

OpenStudy (anonymous):

No problem!!

OpenStudy (anonymous):

if I have to deal with bickering just to know one thing about a asnwer then I should have went to facebook, since it seems people just want to argue over some rule they didn't make up anyways @hero

hero (hero):

There's nothing to argue. 1. OS is about learning 2. Help students understand how to do the problem 3. Don't give answers

OpenStudy (anonymous):

You're being negative @Hero. You broke a rule too

OpenStudy (anonymous):

dude why are you still on that....we got the point, no one asked you for your opinion and we're still not, if you would read you can see that no one asked for the answer jack so get off your high horse

hero (hero):

I'm communicating to users who are unaware of how things work here, how things work here.

hero (hero):

There's nothing negative about that.

hero (hero):

If you have a problem with that, I'll be happy to have a mod explain it to you better.

OpenStudy (anonymous):

do what you feel...just quit crying...

OpenStudy (anonymous):

Yeah. Don't think she appreciates all this argument. Goodbye @Hero and bye( : @mathgirl501

OpenStudy (anonymous):

and for the record...STOP SNITCHING @Hero